# AI.vs.Clinician: Unveiling Intricate Interactions Between AI and Clinicians through an Open-Access Database

### v1

@@ Methods
  • We choose 1500 positive cases and 1500 negative cases considering the demographic diversity, including age, gender, and weight. For positive cases, we also consider balancing the proportion of cases for the three-hour period before onset, the time of onset, and the three-hour period after onset. Note that the total 3000 cases are from 2800 non-repetitive patients. For 100 out of the 2800 patients, we choose three time periods: before, during, and after the onset. For the other 2700 patients, we only include one time period that is either before, during, or after the onset. Finally, we classify these patient cases into five groups, each containing 300 positive cases and 300 negative cases with similar characteristics and distribution.
Table  2 shows the patient cohort population and demographics.
@@
In this step, we perform clinical trials using AI models in our collaborative medical centers. The clinicians will diagnose the patient cases within the well-chosen patient cohort with or without the assistance of different AI models. We build a specialized online diagnosis system for clinicians’ diagnoses and treatments, recording all their operations and timestamps, including the login, logout, diagnosis, treatment, clicked examination items, and all corresponding timestamps. Note that for each patient case, the clinician first performs preliminary diagnosis according to the current fundamental examination data and all the history examination data with or without the probability output of AI models. After the preliminary decision, the clinician can click on the current advanced examination data for more information and make a final decision. The clinicians can choose from five options, severe Sepsis, Sepsis, high suspicion, low suspicion, and non-sepsis, to make a preliminary or final decision, or they can enter the text in the input box directly. In addition, they can provide treatment and drug regimens through a textbox after the preliminary and final diagnosis. All the human-AI interaction data are recorded in this specialized system.
@@
In processing patient data, we strictly adhere to the requirements and specifications of the MIMIC database, conducting thorough data verification and validation. For instance, when using MIMIC-III and MIMIC-IV, as their periods overlap and may contain duplicate patients, we rigorously screen and deduplicate the data to prevent issues of data leakage during AI model training. We select state-of-the-art or state-of-the-practice algorithms for AI models and ensure that the model achieves comparable performance to the reference papers. In the diagnostic phase with clinicians, we construct an online diagnosis system and confirm the correctness and completeness of the data through extensive discussions with numerous clinicians. We strictly adhere to the rules of clinical trials in selecting participating clinicians, setting control groups, and conducting multiple rounds of system usage training. With consent, all actions are recorded by the online system. We verify the completeness and correctness of the data through procedures and manually, such as excluding data where diagnoses are incomplete or irrelevant, to ensure the rationality of the data.
Based on all the records and logs from the online system, the AI.vs.Clinician database is constructed through a series of processes, including integrity checks, consistency checks, and deidentification checks. Specifically, we create a set of rules and unit tests to check all the tables within the database.
@@
Data Access. AI.vs.Clinician database comprises a set of comma-separated value (CSV) files and all related source code. Since the database contains not only the patients’ information from MIMIC databases but also the clinicians’ information from 14 medical centers, users must use the database with caution and respect. The database has been uploaded to PhysioNet [36] platform and is waiting for approval. To access the database, the following steps need to be completed:
@@
Since AI.vs.Clinician database is based on MIMIC database, the users are required to be a credentialed user of MIMIC and complete the required training. Due to the sensitive clinician information, the full diagnosis records and logs cannot be publicly released. However, besides the reorganized database illustrated in this paper, we also provide an original version, which has been performed in the same process of validation and deidentification. The differences are that the original version uses the original table structures from the online diagnosis system and provides both English and Chinese information. After identification, the database is publicly available from PhysioNet (awaiting approval). All the source code for creating patient cohort, AI training and inference, etc., is available from https://github.com/BenchCouncil/AI.vs.Clinician.
@@ Acknowledgements
W.G. conceptualized this study, conceived the experiments, implemented the CoxPHM model according to the referenced paper, and wrote the manuscript. Y.L. conceptualized this study and defined the patient cohort criteria. Z.Y., D.C., W.L., X,L., J.Z., J.X., and H.L. implemented the LSTM models and online diagnosis system and collected and analyzed the data. L.M., N.Y., Y.K., D.L., P.P., W.H., Z.L., J.H., F.H., G.Z., and C.J. recruited clinicians and participated in the experiment. Z.Z., Y.H., and J.Z. conceptualized this study, directed the project, and revised the manuscript. All authors have read and approved the final manuscript.

### v3

@@ Methods
  • We choose 1500 positive cases and 1500 negative cases considering the demographic diversity, including age, gender, and weight. For positive cases, we also consider balancing the proportion of cases for the three-hour period before onset, the time of onset, and the three-hour period after onset. Note that the total 3000 cases are from 2800 non-repetitive patients. For 100 out of the 2800 patients, we choose three time periods: before, during, and after the onset. For the other 2700 patients, we only include one time period that is either before, during, or after the onset. The 3000 cases are organized into five groups, each containing 300 positive cases and 300 negative cases. Table  2 shows the patient cohort population and demographics.
@@
In this step, we perform clinical trials using AI models in our collaborative medical centers. The clinicians will diagnose the patient cases within the well-chosen patient cohort with or without the assistance of different AI models. We build a specialized early warning system for clinicians’ diagnoses and treatments, recording all their operations and timestamps, including the login, logout, diagnosis, treatment, clicked examination items, and all corresponding timestamps. Note that for each patient case, the clinician first performs preliminary diagnosis according to the current fundamental examination data and all the history examination data with or without the probability output of AI models. After the preliminary decision, the clinician can click on the current advanced examination data for more information and make a final decision. The clinicians can choose from five options, severe Sepsis, Sepsis, high suspicion, low suspicion, and non-sepsis, to make a preliminary or final decision, or they can enter the text in the input box directly. In addition, they can provide treatment and drug regimens through a textbox after the preliminary and final diagnosis. All the human-AI interaction data are recorded in this specialized system.
The process of clinician recruitment is as follows:
Step 1: Early warning system on Sepsis implementation, integrating four AI models and a random model.
Step 2: Promotion in multiple medical centers in Guangxi province, China.
Step 3: Conduction of two online training sessions on how to use the early warning system.
Step 4: Clinician recruitment and selection. The inclusion criteria for clinicians are:
(a) possession of a medical license;
(b) having clinical experience in diagnosing and treating sepsis.
Step 5: Early warning system deployment in medical centers where clinicians were recruited.
Step 6: Data collection and analysis.
@@
In processing patient data, we strictly adhere to the requirements and specifications of the MIMIC database, conducting thorough data verification and validation. For instance, when using MIMIC-III and MIMIC-IV, as their periods overlap and may contain duplicate patients, we rigorously screen and deduplicate the data to prevent issues of data leakage during AI model training. We select state-of-the-art or state-of-the-practice algorithms for AI models and ensure that the model achieves comparable performance to the reference papers. In the diagnostic phase with clinicians, we construct an early warning system and confirm the correctness and completeness of the data through extensive discussions with numerous clinicians. We strictly adhere to the rules of clinical trials in selecting participating clinicians, setting control groups, and conducting multiple rounds of system usage training. With consent, all actions are recorded by the early warning system. We verify the completeness and correctness of the data through procedures and manually, such as excluding data where diagnoses are incomplete or irrelevant, to ensure the rationality of the data.
Based on all the records and logs from the early warning system, the AI.vs.Clinician database is constructed through a series of processes, including integrity checks, consistency checks, and deidentification checks. Specifically, we create a set of rules and unit tests to check all the tables within the database.
@@
Data Access. AI.vs.Clinician database comprises a set of comma-separated value (CSV) files and all related source code. Since the database contains not only the patients’ information from MIMIC databases but also the clinicians’ information from 14 medical centers, users must use the database with caution and respect. The database has been uploaded to PhysioNet [36] platform (waiting for approval) under the “Contributor Review" access policy. We also upload the database to the Journal’s online submission system for review. To access the database, the following steps need to be completed:
@@
Since AI.vs.Clinician database is based on MIMIC database, the users are required to be a credentialed user of MIMIC and complete the required training. Due to the sensitive clinician information, the full diagnosis records and logs cannot be publicly released. However, besides the reorganized database illustrated in this paper, we also provide an original version, which has been performed in the same process of validation and deidentification. The differences are that the original version uses the original table structures from the early-warning system and provides both English and Chinese information. After identification, the database is publicly available from PhysioNet (awaiting approval). We also uploaded the database to the journal’s online submission system for review. All the source code for creating patient cohort, AI training and inference, etc., is available from https://github.com/BenchCouncil/AI.vs.Clinician.
@@ Acknowledgements
W.G. conceptualized this study, conceived the experiments, implemented the CoxPHM model according to the referenced paper, and wrote the manuscript. Y.L. conceptualized this study and defined the patient cohort criteria. Z.Y., D.C., W.L., X,L., J.Z., J.X., and H.L. implemented the LSTM models and early warning system and collected and analyzed the data. L.M., N.Y., Y.K., D.L., P.P., W.H., Z.L., J.H., F.H., G.Z., C.J., and T.W. recruited clinicians and participated in the experiment. Z.Z., Y.H., and J.Z. conceptualized this study, directed the project, and revised the manuscript. All authors have read and approved the final manuscript.